# Any-Time Regret-Guaranteed Algorithm for Control of Linear Quadratic Systems

### v2

@@ 6 Discussion and Conclusion
In this section, we discuss the conceptual foundations underlying the findings of this paper to provide a clearer understanding of the challenges involved in algorithm design. One specific discussion concerns the rationale behind the particular choice of perturbation noise, Gaussian $\eta_{t}\sim\mathcal{N}(0,\Gamma_{t})$, with $\|\Gamma_{t}\|=\mathcal{O}\!\big(1/\sqrt{t}\big)$. The following proposition formally justifies why this order of the covariance matrix is desirable for achieving a favorable regret upper bound, and explains how deviations from this order can worsen the regret upper-bound. This choice of scaling has previously been employed in self-tuning–regulator–based algorithms, including CE-based approaches such as simchowitz2020naive and jedra2022minimal.
Proposition 2
@@
In most existing algorithms, due to their specific regret decomposition, the state norm does not directly affect the dominant terms of the regret and usually appears only in $\mathcal{O}(\log T)$ terms. As a result, these works typically do not aim to derive tight high-probability bounds on the state norm trajectory. Their analyses show that stability in the Lyapunov sense is sufficient to control the regret bound, and they often provide only crude bounds on the state norm by arguing that as the length of epochs increases, the effect of policy updates diminishes in the long run. In contrast, we found that carefully studying the boundedness of the state norm is necessary, as it provides strong control over the state trajectory and makes the algorithm, with its anytime regret guarantee, transferable to a wider range of control settings. In our work, we observed that while imposing a strong sequentiality notion can help to explicitly talk about teh state norm bound, the resulting regret bound might not be optimal. This is because strong sequentiality requires additional constraints on the algorithm to keep subsequently generated policies close to each other. We proposed an alternative idea: let the algorithm generate only stabilizing policies, and handle concerns about state explosion due to policy updates using a dwell-time–inspired approach. We noticed that if the focus is purely on minimizing regret, this extreme approach is also inefficient. The optimal regret bound is achieved when the algorithm ensures that two subsequently generated policies are neither too far apart nor too close. This reveals a fundamental trade-off between state norm control and regret upper bounds. Our algorithm lies at the intersection of two broad classes of methods: self-tuning regulator–based algorithms, such as CE-based approaches mania2019certainty; simchowitz2020naive; jedra2022minimal, and OFU-based algorithms, such as abeille2020efficient; cohen2019learning. While our method is in the spirit of OFU, it also incorporates elements of input perturbation methods. This combination allowed us to overcome one of the remaining challenges of computationally efficient OFU-based algorithms: the requirement for an a priori bound on the optimal average expected cost, i.e., $\|P_{*}\|\leq D$. Such a requirement is unnecessary for self-tuning regulator–based algorithms, and by bridging these two approaches, we were able to remove this assumption for OFU-based methods.
@@ Appendix A Discussion of Related Works
The other useful and practical categorization of regret guaranteed algorithms for LQ control, also adopted by jedra2022minimal, is to group existing methods under two main umbrellas: (i) self-tuning regulators and (ii) OFU-based approaches.
@@ Appendix B Preliminaries
B.1 Primal and Dual-SDP formulation of LQR
@@ Appendix D Confidence Ellipsoid Construction
Appendix E Relaxed Primal and Dual SDP
When incorporating the confidence ellipsoid, the primal and dual SDPs (51 and 56) can be relaxed to be used for control design and regret bound analysis. In order to incorporate the estimation error in SDP formulation, we first need the following perturbation lemma. This lemma, a generalized version of Lemma 24 of cohen2019learning, can be used to incorporate un-normalized confidence ellipsoids into SDP formulation.
@@ Appendix E Relaxed Primal and Dual SDP
Now having the perturbation lemma we can write the following lemma to relax the SDPs (51 and 56).
@@ Appendix G Stability Analysis of ARSLO algorithm
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V^{-1}_{t}\preceq\begin{pmatrix}Q&0\\ 0&R\end{pmatrix}$ (90)
@@
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{4}I,$ (91)
@@ G.1 Proof of Theorem 1
where in the inequality (92) we applied the perturbation lemma, Lemma 7. Substituting (93) into (88), yields
@@ H.1 Regret Decomposition
$R_{3}(t)=\sum_{k=1}^{t}\big(w^{\top}_{k}P(\mathcal{C}_{k})w_{k}-\sigma_{\omega}^{2}\|P(\mathcal{C}_{k})\|\big)1_{\mathcal{E}_{k}},$ (139c)
@@ H.2 Proof of Theorem 2
To derive the regret bound, we upper-bound each term in (139a-139f) separately, setting $\beta=1$ to remain consistent with the ARSLO algorithm (Algorithm 1). Throughout the regret analysis, we make use of the norm upper bounds established in Lemma 17.
@@
$\log\frac{\det(V_{t})}{\det(\lambda_{1}I)}\leq 2(n+m)\log\frac{2nt}{\delta}\lesssim(n+m)\log\frac{nt}{\delta}.$
@@
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{4}I\quad\forall t$
@@
Initialize $V_{1}=\lambda I$, $r_{1}=\lambda\bar{\epsilon}^{2}$, $\hat{\Theta}_{1}=\Theta_{0}$, and construct $\mathcal{C}_{1}$.
@@
     Compute $K(\mathcal{C}_{t})$ by (13) and through solving the relaxed primal SDP problem (12)
@@
  Construct $X_{t}$ and $Z_{t}$ and compute $V_{t}$ using regularization parameter $\lambda=\mathcal{G}\log\frac{2nt}{\delta}$
@@
$P_{*}-P(\mathcal{C}_{t})$ $\preceq\frac{16\|P(\mathcal{C}_{t})\|^{2}\mu_{t}}{\alpha_{0}^{2}}\|P(\mathcal{C}_{t})\|_{*}\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V^{-1}_{t}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}\|I$
$=4\kappa_{t}^{4}\mu_{t}\|P(\mathcal{C}_{t})\|_{*}\|\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}^{\top}V^{-1}_{t}\begin{pmatrix}I\\ K(\mathcal{C}_{t})\end{pmatrix}\|I.$ (174)
@@
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}\|V_{t}^{-1}\|I\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I$ (175)
@@
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I$
@@
L.1 Proof of Theorem 6
Proof Recall that $\Theta_{*}$ belongs to the confidence ellipsoid $\mathcal{C}_{t}^{w}(\delta)$ with probability $1-\delta$, which implies
$\operatorname{Tr}\Big((\Theta_{*}-\hat{\Theta}^{w}_{t})^{\top}\bar{V}_{t}(\Theta_{*}-\hat{\Theta}^{w}_{t})\Big)\leq r^{w}_{t}.$
It then follows that
$\|\Theta_{*}-\hat{\Theta}^{w}_{t}\|_{F}^{2}\leq\frac{r_{t}^{w}}{\lambda_{\min}(\bar{V}_{t})}.$
Hence, a sufficient condition to guarantee
$\|\Theta_{*}-\hat{\Theta}^{w}_{t}\|_{F}\leq\varepsilon$
where $\varepsilon$ can be either $\bar{\epsilon}(\bar{\kappa}_{1})$ or $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$ is
$\sqrt{\frac{r_{t}^{w}}{\lambda_{\min}(\bar{V}_{t})}}\leq\varepsilon,$ (190)
which completes the proof.
@@
L.2 Proof of Theorem 7
Proof Let $t_{w}$ denote the duration of the warm-up phase which is determined by termination condition of the warm-up phase (190):
@@
Appendix M Proof of Proposition 2

### v3

@@ 6 Discussion and Conclusion
In most existing algorithms, due to their specific regret decomposition, the state norm does not directly affect the dominant terms of the regret and usually appears only in $\mathcal{O}(\log T)$ terms. As a result, these works typically do not aim to derive tight high-probability bounds on the state norm trajectory. Their analyses show that stability in the Lyapunov sense is sufficient to control the regret bound, and they often provide only crude bounds on the state norm by arguing that as the length of epochs increases, the effect of policy updates diminishes in the long run. In contrast, we found that carefully studying the boundedness of the state norm is necessary, as it provides strong control over the state trajectory and makes the algorithm, with its anytime regret guarantee, transferable to a wider range of control settings. In our work, we observed that while imposing a strong sequentiality notion can help to explicitly talk about the state norm bound, the resulting regret bound might not be optimal. This is because strong sequentiality requires additional constraints on the algorithm to keep subsequently generated policies close to each other. We proposed an alternative idea: let the algorithm generate only stabilizing policies, and handle concerns about state explosion due to policy updates using a dwell-time–inspired approach. We noticed that if the focus is purely on minimizing regret, this extreme approach is also inefficient. The optimal regret bound is achieved when the algorithm ensures that two subsequently generated policies are neither too far apart nor too close. This reveals a fundamental trade-off between state norm control and regret upper bounds. Our algorithm lies at the intersection of two broad classes of methods: self-tuning regulator–based algorithms, such as CE-based approaches mania2019certainty; simchowitz2020naive; jedra2022minimal, and OFU-based algorithms, such as abeille2020efficient; cohen2019learning. While our method is in the spirit of OFU, it also incorporates elements of input perturbation methods. This combination allowed us to overcome one of the remaining challenges of computationally efficient OFU-based algorithms: the requirement for an a priori bound on the optimal average expected cost, i.e., $\|P_{*}\|\leq D$. Such a requirement is unnecessary for self-tuning regulator–based algorithms, and by bridging these two approaches, we were able to remove this assumption for OFU-based methods. Finally, although the techniques developed in this work are derived for LQ control with regret guarantees, they may be extended to more general learning-based nonlinear control schemes. In such settings, where policy updates occur over time, the resulting closed-loop dynamics can naturally be interpreted as switched systems. This perspective suggests a principled pathway for relating closed-loop stability, performance, and the exploration–exploitation trade-off through tools from switched-systems theory, such as dwell-time.
@@ Appendix A Discussion of Related Works
One useful and practical categorization of regret guaranteed algorithms for LQ control, also adopted by jedra2022minimal, is to group existing methods under two main umbrellas: (i) self-tuning regulators and (ii) OFU-based approaches.
@@ Appendix B Preliminaries
B.1 Primal and Dual SDP Formulation of LQR
@@ Appendix D Confidence Ellipsoid Construction
Appendix E Relaxed Primal and Dual SDPs
When incorporating the confidence ellipsoid, the primal and dual SDPs (53 and 58) can be relaxed to be used for control design and regret bound analysis. In order to incorporate the estimation error in SDP formulation, we first need the following perturbation lemma. This lemma, a generalized version of Lemma 24 of cohen2019learning, can be used to incorporate un-normalized confidence ellipsoids into SDP formulation.
@@ Appendix E Relaxed Primal and Dual SDPs
Now having the perturbation lemma we can write the following lemma to relax the SDPs (53 and 58).
@@ Appendix G Stability Analysis of ARSLO Algorithm
$\mu_{t}\|P(\mathcal{C}_{t})\|V^{-1}_{t}\preceq\begin{pmatrix}Q&0\\ 0&R\end{pmatrix}$ (92)
@@
$\mu_{t}\|P(\mathcal{C}_{t})\|V_{t}^{-1}\preceq\frac{\alpha_{0}}{4}I,$ (93)
@@ G.1 Proof of Theorem 1
where $\hat{\Theta}_{t}=(\hat{A}_{t},\,\hat{B}_{t})^{\top}$ and in the inequality (94) we applied the perturbation lemma, Lemma 7. Substituting (95) into (90), yields
@@ H.1 Regret Decomposition
$R_{3}(t)=\sum_{k=1}^{t}\big(w^{\top}_{k}P(\mathcal{C}_{k})w_{k}-\sigma_{\omega}^{2}\|P(\mathcal{C}_{k})\|_{*}\big)1_{\mathcal{E}_{k}},$ (141c)
@@ H.2 Proof of Theorem 2
To derive the regret bound, we upper-bound each term in (141a-141f) separately, setting $\beta=1$ to remain consistent with the ARSLO algorithm (Algorithm 1). Throughout the regret analysis, we make use of the norm upper bounds established in Lemma 17.
@@
$\log\frac{\det(V_{t})}{\det(\lambda I)}\leq 2(n+m)\log\frac{2nt}{\delta}\lesssim(n+m)\log\frac{nt}{\delta}.$
@@
$\mu_{t}\|P(\mathcal{C}_{t})\|V_{t}^{-1}\preceq\frac{\alpha_{0}}{4}I\quad\forall t$
@@
Initialize $V_{1}=\lambda I$, $r_{1}=\lambda\underline{\epsilon}^{2}(\bar{\kappa}_{1},\bar{\rho})$, $\hat{\Theta}_{1}=\Theta_{0}$, and construct $\mathcal{C}_{1}$.
@@
     Solve the relaxed primal SDP problem (12) and compute $K(\mathcal{C}_{t})$ by (13)
@@
  Construct $X_{t}$, $Z_{t}$ and update $V_{t}$
@@
$\mu_{t}\|P(\mathcal{C}_{t})\|\|V_{t}^{-1}\|I\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I$ (177)
@@
$\mu_{t}\|P(\mathcal{C}_{t})\|V_{t}^{-1}\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I$
@@
L.1 Proof of Theorem 6
Proof Let $t_{w}$ denote the duration of the warm-up phase which is determined by termination condition of the warm-up phase $\epsilon_{w}(t)\leq\varepsilon$:
@@
Appendix M Rationale for the Perturbation Noise Scaling
Throughout this paper, for both proposed algorithms, we choose Gaussian perturbation noise $\eta_{t}\sim\mathcal{N}(0,\Gamma_{t})$, with $\|\Gamma_{t}\|=\mathcal{O}\!\big(1/\sqrt{t}\big)$. However, we have not yet discussed the rationale behind this particular choice. The following proposition formally justifies why this order of the covariance matrix is desirable for achieving a favorable regret upper bound, and explains how deviations from this scaling can worsen the regret bound. This choice of scaling has previously been employed in self-tuning–regulator–based algorithms, including CE-based approaches such as simchowitz2020naive and jedra2022minimal.
Proposition 2